# Unconstrained Open Vocabulary Image Classification: Zero-Shot Transfer from Text to Image via CLIP Inversion

### v2

@@ Abstract
We introduce NOVIC, an innovative uNconstrained Open Vocabulary Image Classifier that uses an autoregressive transformer to generatively output classification labels as language. Leveraging the extensive knowledge of CLIP models, NOVIC harnesses the embedding space to enable zero-shot transfer from pure text to images. Traditional CLIP models, despite their ability for open vocabulary classification, require an exhaustive prompt of potential class labels, restricting their application to images of known content or context. To address this, we propose an “object decoder” model that is trained on a large-scale 92M-target dataset of templated object noun sets and LLM-generated captions to always output the object noun in question. This effectively inverts the CLIP text encoder and allows textual object labels to be generated directly from image-derived embedding vectors, without requiring any a priori knowledge of the potential content of an image. The trained decoders are tested on a mix of manually and web-curated datasets, as well as standard image classification benchmarks, and achieve fine-grained prompt-free prediction scores of up to 87.5%, a strong result considering the model must work for any conceivable image and without any contextual clues.
@@ 1 Introduction
In summary, our main contributions [^1] All code and datasets will be made publicly available. are as follows: (i) we propose a novel open vocabulary object decoder model that is trained on text only and can zero-shot classify arbitrary images in real-time without any candidates or prompts, (ii) we develop an automated pipeline to construct a comprehensive English object noun dictionary and use a multiset prompting scheme combined with an LLM to construct large-scale synthetic caption-object datasets, (iii) we curate three new image sets intended for evaluation of open vocabulary classification and provide human and multimodal LLM annotations, and (iv) we show that our method scales in performance with the underlying CLIP model, and is able to provide accurate yet very fine-grained classifications while achieving strong in-the-wild prediction scores up to 87.5%.
@@ 2 Related Work
A recent line of research on zero-shot image captioning or retrieval with VLMs proposes text-only training to eliminate the need for curating expensive image datasets [39]. As text and images are encoded into disjoint regions of the VLM embedding space, additional mechanisms to bridge the modality gap have been proposed, like injecting noise into the text embeddings [45, 18, 16] or projecting the image embeddings [36]. These methods face two significant disadvantages—first, captions often fail to capture key image aspects or identify core objects, and second, they are biased towards the limited set of object classes contained in their training dataset. NOVIC adopts text-only training, but focuses on predicting truly open vocabulary object labels.
@@ 3.1 Object Decoder Model
The architecture of the object decoder model is shown in Fig. 3. The model is built around an autoregressive decoder-only language transformer [57] that has been configured to use learned token embeddings, weight tying [47], learned positional encodings, pre-layer normalization [64], GELU activation functions [22], beam search, custom weight initialization (cf. App. A), a contractive inner feed-forward dimension, and no biases in the linear and layer normalization layers, including the final token logits linear layer. Dropout is applied after adding the positional encoding and within the transformer layers themselves. The input embedding vector is supplied to the transformer in the form of $P$ prefix token vectors, i.e. $P$ dynamically calculated vectors of dimension $H$ (hidden dimension). The prefix tokens are computed from the input embedding using a learned linear projection that has $PH$ outputs and no biases, as the immediately following positional encoding makes biases redundant. The tokenized target object nouns do not require a start token, as the first output token is predicted at the sequence location of the last prefix token. Appropriate sequence masking is also used to ensure that only sequence locations up to the first output <eos> contribute to the loss. The internal attention mask of the transformer is causal, with the exception that the $P$ prefix tokens can attend to each other without restriction, allowing them to share information freely.
@@ 3.2 Training Dataset for Text-Only Supervision
An exhaustive dictionary of simple and compound object nouns, referred to as the object noun dictionary, was created based on the WordNet dictionary [42], GNU Collaborative International Dictionary of English [5], and the categories used by various image datasets (cf. App. B). [^2] A best effort was made to keep the object noun dictionary as accurate and complete as possible given the available dictionary data. A missing word would be indicative of a possible improvement to the object noun dictionary rather than a weakness of the object decoder training scheme. Noun entries were cleaned, collected and merged as appropriate to account for word duplication, plurality, inflection and alternate spellings, and filtered down to only include nouns that correspond to suitable object classifications (e.g. no abstract nouns). This resulted in 42 919 unique case-normalized object nouns corresponding to 96 526 noun variants. Each object noun was then assigned a log-scale frequency value based on its observed frequency in the English language as per the n-gram frequencies in the Google Web Trillion Word Corpus [44, 4]. Object nouns that are categories in a curated, comprehensive collection of image datasets were given appropriate frequency boosts. The log-scale frequencies were scaled and rounded so that very rare nouns saturate at a value of 1, and very frequent nouns have values $\geq$12. Aside from being used to adjust relative sampling frequencies during training, the resulting frequencies allowed the construction of object noun dictionary subsets by filtering out nouns less or equal to a particular frequency threshold (FT), e.g. FT2.333FT0, FT2, FT6, and FT9 have 42 919, 11 897, 5 899, and 2 919 unique object nouns, respectively. FT0 is the complete object noun dictionary.
@@ 3.3 Evaluation Datasets for Zero-Shot Recognition
As NOVIC is trained without providing explicit categories, we wish to evaluate it also on datasets that do not specify categories. No suitable such dataset is known however, as such a dataset would need to have annotations for every possible correct object noun in the English language for every visible entity in every part of each image. To address this, three open vocabulary image datasets were curated as part of this work (cf. App. E), and individually annotated by both human and multimodal LLM annotators (cf. App. F). The annotations specify whether each classification is correct, close, or incorrect, and for the human annotations, whether it relates to a primary or secondary element of the image. The three new datasets are (i) World: 272 images of which the grand majority are originally sourced (have never been on the internet) from 10 countries by 12 people, with an active focus on covering as wide and varied concepts as possible, including unusual, deceptive and/or indirect representations of objects, (ii) Wiki: 1 000 Wikipedia lead images sampled from a scraped pool of 18K, (iii) Val3K: 3 000 images from the ImageNet-1K validation set, sampled uniformly across the classes (LLM annotations only). The suffixes -H and -L are used to clarify whether human or LLM annotations are in use, e.g. Wiki-H. A total of 16.5K human and 112K LLM (OpenAI GPT-4o) annotations were made across the three datasets. This was at the limit of feasibility in terms of number of samples, person-hours and cost.
@@ 4.1 Zero-Shot Classification Performance
Fashion MNIST [63] 68.41 72.59 70.70 70.64 65.74
CIFAR-10 [31] 92.33 90.75 90.56 90.79 91.00
@@
Food-101 [3] 91.55 85.25 83.70 83.61 70.49
Tiny ImageNet [33] 59.68 56.98 57.36 55.83 53.78
Imagenette [12] 99.59 99.44 99.29 99.32 99.23
Imagewoof [12] 93.28 93.07 92.09 92.20 90.15
ImageNet-1K [49] 76.06 46.76 69.50 68.11 60.56
Table 1: Performance on image classification benchmarks (mean of 3). FT9 has an outlier ImageNet-1K score as it has not seen enough object nouns to cover all 1000 classes.
Object nouns FT9 FT6 FT2 FT0
ImageNet-1K (3K) 46.52 69.47 67.92 60.72
@@
Tab. 1 shows the zero-shot performance of NOVIC on various image classification benchmarks. The performance of the underlying CLIP model is also shown as a measure of how completely the vast knowledge of the CLIP model has been captured by the object decoder and converted into generative textual understanding. The decoder relies on the organization of the embedding space provided by the CLIP model, and so naturally is not expected to outperform it in this category-prompted setting. Overall, NOVIC shows competitive top-1 accuracies, with some ‘distillation’ losses being observed for tendentially finer-grained classification datasets and lower frequency thresholds. Classification benchmarks, however, do not tell the whole story. The act of restricting the object decoder to a fixed vocabulary forces it to choose object nouns it does not even consider to be best. In Tab. 2, we evaluate the top-1 accuracy on Val3K using the original ImageNet-1K annotations, and compare this to the percent of open vocabulary classifications on the very same images that are correct as per Val3K-L. Prediction scores that include points for close predictions are also shown. The results clearly demonstrate that the object decoder’s knowledge is more accurately assessed when used, as intended, for unconstrained open vocabulary classification.
@@
As discussed in Sec. 2, no directly equivalent state-of-the-art model is known that performs real-time prompt-free open vocabulary image classification. The most suitable models for comparison are image tagging models like Tag2Text [24] and the Recognize Anything Model (RAM) [72], as they are designed to be able to provide text labels for images of arbitrary content. We evaluate the largest model variants (14M) on Wiki-H by taking the highest scoring prediction for each image, and compare this to our results in Tab. 3. While the tagging models obtain a high numeric prediction score, a closer look reveals this is due to a severe lack of diversity and specificity (cf. Figs. 4 and H). The produced tags are tendentially coarse and repetitive, with around a third of all images receiving one of only 20 generic top labels, often missing the main point of images. Even when provided with the full object noun dictionary, the total diversity of labels stays below 400, with over 80% of predictions still being made from the small core RAM vocabulary. This is in contrast to our model, which has a much flatter prediction distribution, diversities up to 850 (cf. Tab. 3), and a high prediction score despite generating fine-grained classifications. The tags produced by Tag2Text and RAM also tend to be uninformative—when applied to the images of the Imagewoof [12] validation set for example, both models simply output dog 88% of the time, and never a single dog breed. Our FT0 model, on the other hand, predicts over 90 different dog breeds, with much of the weight covering variations of the actual ground truth classes. When RAM is evaluated on ImageNet-1K by only providing it the 1000 class names to choose from, it only scores 38.3% (25.4% for non-core vocabulary), which is far less than the 75.54% of the ViT-L CLIP model it was trained on, and far less than the 60–70% nominally scored by NOVIC.
@@
Baseline Text-only Training Seen Objects Used Objects RAM Vocab (%) Prediction Score (%)
Tag2Text [24] 3 429 397 89.0 81.35
RAM [72] 4 585 348 100.0 80.73
RAM (FT2) 11 897 391 83.4 76.94
RAM (FT0) 42 919 399 81.7 77.62
NOVIC (FT9) 2 919 693 64.0 72.35
NOVIC (FT6) 5 899 794 49.3 72.70
NOVIC (FT2) 11 897 846 39.7 72.53
NOVIC (FT0) 42 919 852 37.5 72.34
@@
Table 3: Open vocabulary Wiki-H results in comparison to image tagging models (best of 3). RAM and Tag2Text achieve artificially high scores by frequently outputting uninformative generic tags like man, tree, grass, plate, and not for instance what is on the grass or plate, or what type of tree it is, or what the man is using.
@@ 4.2 Ablation Studies
The effect of the object noun dictionary frequency threshold (cf. Sec. 3.2) on decoder classification performance is shown in Tab. 4. The prediction scores grow as the training dataset size is increased, until a limit is reached between FT2 and FT0 where the exploding diversity of seen object nouns saturates the amount of order the CLIP model can provide in the embedding space. In Tab. 4, we systematically analyze the contribution of each component of the training set. Starting with $\mathcal{M}_{1}$ and incrementally building the nominal training set (i.e. $\mathcal{M}_{1}$ + $\mathcal{M}_{2}$ + $\mathcal{M}_{3}$ + LLM captions), we observe consistent gains. Immediate losses are seen if the LLM captions are removed, or if the lower-order $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ multisets are removed. Further tests with variations in training noise configuration also validate that very large [^7] Gauss 3.25 in the table refers to elementwise Gaussian noise with a standard deviation so high that the expected norm of the noise vector that is added to the unit-norm embedding is 3.25, i.e. many times its magnitude. amounts of Gaussian noise are required in order to produce models that can generalize across the modality gap, and that instead using uniform angle noise on 15% of the samples produces the best seen configuration.
CLIP model.
Language generalization.
During training, the object decoder develops an implicit understanding of language and its relationship to object concepts, demonstrating qualitative evidence that it can occasionally use this knowledge to generalize beyond the dataset. For instance, the unseen noun New Zealand willow was predicted for a scenic photo of a willow tree that was indeed taken in New Zealand. The object decoder also inherits CLIP’s slight tendency to be able to ‘read’, as evidenced for example by a generic blue plushie being correctly classified by multiple decoders as genus staphylococcus, but not anymore if the small visible text Staphylokokke is removed, falling back to stuffed toy. In another case, a sign reading Aral is classified as ara (a macaw), indicating that the decoder can associate certain features in the embedding space with certain spellings.
@@
Table 4: Dataset, training noise, and CLIP model ablations (mean of 3). IN1K is the ImageNet-1K classification performance of the trained decoders, and the asterisks correspond to the nominal ablation baseline. See App. I for more ablation results.

### v3

@@ Abstract
We introduce NOVIC, an innovative real-time uNconstrained Open Vocabulary Image Classifier that uses an autoregressive transformer to generatively output classification labels as language. Leveraging the extensive knowledge of CLIP models, NOVIC harnesses the embedding space to enable zero-shot transfer from pure text to images. Traditional CLIP models, despite their ability for open vocabulary classification, require an exhaustive prompt of potential class labels, restricting their application to images of known content or context. To address this, we propose an “object decoder” model that is trained on a large-scale 92M-target dataset of templated object noun sets and LLM-generated captions to always output the object noun in question. This effectively inverts the CLIP text encoder and allows textual object labels from essentially the entire English language to be generated directly from image-derived embedding vectors, without requiring any a priori knowledge of the potential content of an image, and without any label biases. The trained decoders are tested on a mix of manually and web-curated datasets, as well as standard image classification benchmarks, and achieve fine-grained prompt-free prediction scores of up to 87.5%, a strong result considering the model must work for any conceivable image and without any contextual clues. [^1] The authors gratefully acknowledge support from the DFG (TRR 169 – CML) and European Commission (TRAIL).
@@ 1 Introduction
In summary, our main contributions [^2] All code [3, 2] (GitHub) and datasets [4, 5] are publicly available. are as follows: (i) we propose a novel open vocabulary object decoder model that is trained on text only and can zero-shot classify arbitrary images in real-time without any candidates or prompts, (ii) we develop an automated pipeline to construct a comprehensive English object noun dictionary and use a multiset prompting scheme combined with an LLM to construct large-scale synthetic caption-object datasets, (iii) we curate three new image sets intended for evaluation of open vocabulary classification and provide human and multimodal LLM annotations, and (iv) we show that our method scales in performance with the underlying CLIP model, and is able to provide accurate yet very fine-grained classifications while achieving strong in-the-wild prediction scores up to 87.5%.
@@ 2 Related Work
A recent line of research on zero-shot image captioning or retrieval with VLMs proposes text-only training to eliminate the need for curating expensive image datasets. As text and images are encoded into disjoint regions of the VLM embedding space, additional mechanisms to bridge the modality gap have been proposed, like injecting noise into the text embeddings [50, 22, 20] or projecting the image embeddings [41]. These methods face two significant disadvantages—first, captions often fail to capture key image aspects or identify core objects, and second, they are biased towards the limited set of object classes contained in their training dataset. NOVIC adopts text-only training, but focuses on predicting truly open vocabulary object labels.
@@ 3.1 Object Decoder Model
The architecture of the object decoder model is shown in Fig. 3. The model is built around an autoregressive decoder-only language transformer [62] that has been configured to use learned token embeddings, weight tying [52], learned positional embeddings, pre-layer normalization [68], GELU activation functions [26], beam search, custom weight initialization (cf. App. A), a contractive inner feed-forward dimension, and no biases in the linear and layer normalization layers, including the final token logits linear layer. Dropout is applied after the positional embedding and within the transformer layers. The input embedding vector is supplied to the transformer in the form of $P$ prefix token vectors, i.e. $P$ dynamically calculated vectors of dimension $H$ (hidden dimension). The prefix tokens are computed from the input embedding using a learned linear projection that has $PH$ outputs and no biases, as the immediately following positional embedding makes biases redundant. The tokenized target object nouns do not require a start token, as the first output token is predicted at the sequence location of the last prefix token. Appropriate sequence masking is also used to ensure that only sequence locations up to the first output <eos> contribute to the loss. The internal attention mask of the transformer is causal, with the exception that the $P$ prefix tokens can attend to each other without restriction, allowing them to share information freely.
@@ 3.2 Training Dataset for Text-Only Supervision
An exhaustive dictionary of simple and compound object nouns, referred to as the object noun dictionary, was created based on the WordNet dictionary [47], GNU Collaborative International Dictionary of English [9], and the categories used by various image datasets (cf. App. B). Noun entries were cleaned, collected and merged as appropriate to account for word duplication, plurality, inflection and alternate spellings, and filtered down to only include nouns that correspond to suitable object classifications (e.g. no abstract nouns). This resulted in 42 919 unique case-normalized object nouns corresponding to 96 526 noun variants. Each object noun was then assigned a log-scale frequency value based on its observed frequency in the English language as per the n-gram frequencies in the Google Web Trillion Word Corpus [49, 8]. Object nouns that are categories in a curated, comprehensive collection of image datasets were given appropriate frequency boosts. The log-scale frequencies were scaled and rounded so that very rare nouns saturate at a value of 1, and very frequent nouns have values $\geq$12. Aside from being used to adjust relative sampling frequencies during training, the resulting frequencies allowed the construction of object noun dictionary subsets by filtering out nouns less or equal to a particular frequency threshold (FT), e.g. FT2.333FT0, FT2, FT6, and FT9 have 42 919, 11 897, 5 899, and 2 919 unique object nouns, respectively. FT0 is the complete object noun dictionary.
@@ 3.3 Evaluation Datasets for Zero-Shot Recognition
As NOVIC is trained without providing explicit categories, we wish to evaluate it also on datasets that do not specify categories. No suitable such dataset is known however, as such a dataset would need to have annotations for every possible correct object noun in the English language for every visible entity in every part of each image. To address this, three open vocabulary image datasets were curated as part of this work (cf. App. E), and individually annotated by both human and multimodal LLM annotators (cf. App. F) for the object nouns that were predicted by the trained models. The annotations specify whether each classification is correct, close, or incorrect, and for the human annotations, whether it relates to a primary or secondary element of the image. The three new datasets are (i) World: 272 images of which the grand majority are originally sourced (have never been on the internet) from 10 countries by 12 people, with an active focus on covering as wide and varied concepts as possible, including unusual, deceptive and/or indirect representations of objects, (ii) Wiki: 1 000 Wikipedia lead images sampled from a scraped pool of 18K, (iii) Val3K: 3 000 images from the ImageNet-1K validation set, sampled uniformly across the classes (LLM annotations only). The suffixes -H and -L are used to clarify whether human or LLM annotations are in use, e.g. Wiki-H. A total of 17.4K human and 112K LLM (OpenAI GPT-4o) annotations were made across the three datasets. This was at the limit of feasibility in terms of number of samples, person-hours and cost.
@@ 4.1 Zero-Shot Classification Performance
ImageNet-1K [54] 75.91 46.76 69.50 68.11 60.56
Tiny ImageNet [38] 59.77 54.53 57.36 55.83 53.78
Imagenette [16] 99.59 99.44 99.27 99.32 99.23
Imagewoof [16] 93.28 93.07 92.07 92.20 90.15
ImageNet-A [27] 45.05 35.33 42.43 41.42 38.23
ImageNet-R [27] 90.24 74.46 88.01 86.62 83.43
Food-101 [7] 91.55 85.25 83.24 82.73 68.80
CIFAR-10 [36] 92.33 91.08 90.53 90.77 90.92
@@
Table 1: Performance on image classification benchmarks (mean of 3). FT9 has disadvantaged scores on some datasets due to the seen object nouns not covering all the classes (cf. App. J).
Annotations FT9 FT6 FT2 FT0
ImageNet-1K 46.52 69.47 67.92 60.72
@@
Tab. 1 shows the zero-shot performance of NOVIC on various image classification benchmarks. The performance of the underlying CLIP model is also shown as a measure of how completely the vast knowledge of the CLIP model has been captured by the object decoder and converted into generative textual understanding. The decoder relies on the organization of the embedding space provided by the CLIP model, and so naturally is not expected to outperform it in this category-prompted setting. Overall, NOVIC shows competitive top-1 accuracies, with some ‘distillation’ losses being observed for tendentially finer-grained classification datasets and lower frequency thresholds. Classification benchmarks, however, do not tell the whole story. The act of restricting the object decoder to a fixed vocabulary forces it to choose object nouns it does not even consider to be best. In Tab. 2, we evaluate the top-1 accuracy on Val3K using the original ImageNet-1K annotations, and compare this to the percent of open vocabulary classifications on the very same images that are correct as per Val3K-L. Prediction scores that include points for close predictions are also shown. The results demonstrate that the object decoder’s knowledge is more accurately assessed when used for unconstrained open vocabulary classification.
@@
As discussed in Sec. 2, no directly equivalent state-of-the-art model is known that performs real-time prompt-free open vocabulary image classification. The most suitable models for comparison are image tagging models like Tag2Text [29] and the Recognize Anything Model (RAM) [76], as they are designed to be able to provide text labels for images of arbitrary content. We evaluate the largest model variants by taking the highest scoring prediction for each image, and compare this to NOVIC in Tabs. 3 and K. While the tagging models at first glance numerically seem to obtain a reasonable Prediction Score, the comparatively low Primary Scores indicate that the models, unlike NOVIC, make a large portion of their score from secondary or close predictions, as opposed to correct primary ones (a deficiency clearly illustrated in Tab. 6). Also, when accounting for prediction specificity, the resulting Overall Score clearly demonstrates the superiority of NOVIC by up to 19% absolute. A closer look reveals that the low scores of the tagging models are due to a severe lack of diversity and specificity (cf. Fig. 4 and qualitative results in App. H). The produced tags are tendentially coarse and repetitive, with around a third of all images receiving one of only 20 generic top labels, often missing the main point of images. Even when provided with the full object noun dictionary, the total diversity of labels stays below 400, with over 80% of predictions still being made from the small core RAM vocabulary. This is in contrast to our model, which has a much flatter prediction distribution, diversities up to 934 (cf. Tab. 3), and a remarkably high prediction score despite generating fine-grained classifications. The tags produced by Tag2Text and RAM also tend to be uninformative—when applied to the images of the Imagewoof [16] validation set for example, both models simply output dog 88% of the time, and never a single dog breed. Our FT0 model, on the other hand, predicts over 90 different dog breeds, with much of the weight covering variations of the actual ground truth classes. When RAM is evaluated on ImageNet-1K by only providing it the 1000 class names to choose from, it only scores 38.3% (25.4% for non-core vocabulary), which is far less than the 75.54% of the ViT-L CLIP model it was trained on, and far less than the 60–70% nominally scored by NOVIC.
@@
Model Seen Objects Used Objects RAM Vocab Primary Score Prediction Score Overall Score
Tag2Text [29] 3 429 397 89.0 69.15 81.35 62.61
RAM [76] 4 585 348 100.0 65.05 80.73 59.98
RAM (FT2) 11 897 391 83.4 62.30 77.62 60.58
RAM (FT0) 42 919 399 81.7 61.70 76.94 60.25
NOVIC (FT9) 2 919 693 64.0 68.55 72.35 69.08
NOVIC (FT6) 5 899 794 49.3 69.90 72.70 71.07
NOVIC (FT2) 11 897 846 39.7 68.65 72.53 71.88
NOVIC (FT0) 42 919 852 37.5 68.70 72.34 71.31
NOVIC (FT2) 11 897 893 36.3 77.05 80.13 79.02
NOVIC (FT0) 42 919 934 28.7 77.10 79.18 78.21
@@
Table 3: Open vocabulary Wiki-H results in comparison to baselines (best of 3). Tag2Text and RAM achieve artificially high scores by frequently outputting uninformative generic tags like man, tree, grass, plate, and not for instance what is on the grass or plate, or what type of tree it is, or what the man is using. Scores are in %. RAM Vocab is the percent of samples whose predictions are in the core RAM vocabulary. Primary Score is the Prediction Score when rewarding only primary predictions. Overall Score is the Prediction Score with $\times$0.5 applied to the scores of coarse predictions. See App. K. Using DFN-5B H/14-378 CLIP model.
@@ 4.2 Ablation Studies
The effect of the object noun dictionary frequency threshold (cf. Sec. 3.2) on decoder classification performance is shown in Tab. 4. The prediction scores grow as the training dataset size is increased, until a limit is reached between FT2 and FT0 where the exploding diversity of seen object nouns saturates the amount of order the CLIP model can provide in the embedding space. In Tab. 4, we systematically analyze the contribution of each component of the training set. Starting with $\mathcal{M}_{1}$ and incrementally building the nominal training set (i.e. $\mathcal{M}_{1}$ + $\mathcal{M}_{2}$ + $\mathcal{M}_{3}$ + LLM captions), we observe consistent gains. Immediate losses are seen if the LLM captions are removed, or if the lower-order $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ multisets are removed. Further tests with variations in training noise configuration also validate that very large amounts of Gaussian noise are required in order to produce models that can generalize across the modality gap, [^7] Gauss 3.25 in the table refers to elementwise Gaussian noise with a standard deviation so high that the expected norm of the noise vector that is added to the unit-norm embedding is 3.25, i.e. many times its magnitude. and that instead using uniform angle noise on 15% of the samples produces the best seen configuration.
@@
Table 4: Dataset, training noise, and CLIP model ablations (mean of 3). IN1K is the ImageNet-1K classification performance of the trained decoders, and the asterisks correspond to the nominal ablation baseline. See App. I for more ablation results. Some Wiki-H scores are unavailable due to the high cost of manual annotation.
CLIP model.
Language generalization.
During training, the object decoder develops an implicit understanding of language and its relationship to object concepts, demonstrating qualitative evidence that it can occasionally use this knowledge to generalize beyond the dataset. For instance, the unseen noun New Zealand willow was predicted for a scenic photo of a willow tree that was indeed taken in New Zealand. The object decoder also inherits CLIP’s slight tendency to be able to ‘read’, as evidenced for example by a generic blue plushie being correctly classified by multiple decoders as genus staphylococcus, but not anymore if the small visible text Staphylokokke is removed, falling back to stuffed toy. In another case, a sign reading Aral is classified as ara (a macaw), indicating that the decoder can associate certain features in the embedding space with certain spellings. As another form of language generalization, the object decoder can easily be trained to output high-performing predictions in any chosen language—even low resource languages for which insufficient image-text data exists to train a CLIP model. By simply applying dictionary or machine translation to the target object nouns in the $\mathcal{M}_{m}$ multisets, a predominantly English CLIP model can be used to directly train a strong object decoder in any translatable language.